# Pre-processing and Compression: Understanding Hidden Representation Refinement Across Imaging Domains via Intrinsic Dimension

### v3

@@ Introduction
In this work, we first explore how layer-by-layer representation ID progression differs for models trained on datasets beyond ImageNet, including both three other natural image datasets and seven medical image datasets for a wide range of diagnostic tasks. We find that these curves have typically different shapes depending on if the model was trained on natural or medical images. In particular, the depth at which the ID peaks at some $d_{\mathrm{repr}}^{{\mathrm{max}}}$ is noticeably deeper into the model on average for natural image models, which we attribute to the task-relevant features of medical images typically requiring less abstract representations to capture, and so can be “pre-processed” for compression earlier in the network. Next, we explore the relationship between $d_{\mathrm{repr}}^{{\mathrm{max}}}$ and the intrinsic dimension of the raw data in its native space, $d_{\mathrm{data}}$ [19], and discover a new result: there is a strong correlation between the two across all datasets and models, showing that the maximum intrinsic complexity of learned representations in a model is guided by the intrinsic complexity of the model’s training data.
@@ 1 Experimental Methods
Neural networks work by mapping high-dimensional input data (like images) to significantly lower-dimensional manifolds of learned representations [6, 1] which describe features that generalize to new data, made possible by learning the lower-dimensional manifold structure of such datasets in their native space [4]. The intrinsic dimension of such a manifold can be computed via a relatively simple estimator that utilizes maximum likelihood estimation (MLE) [14, 16] given the high dimensional dataset (in our case, of representations) which we use here following prior work [11, 19] via the implementation of https://github.com/mazurowski-lab/intrinsic-properties.
@@ 2 Experiments and Results
A manifold hypothesis-related idea is that the learned representation manifolds of a network may be shaped by the properties and complexity of the dataset manifold in its native (pixel) space. Prior work [11] studied the relationship between the intrinsic dimension of final hidden layer representations for classification networks, $d_{\mathrm{repr}}^{L-1}$, with the intrinsic dimension of the dataset itself, $d_{\mathrm{data}}$, and theoretically and empirically showed that $d_{\mathrm{repr}}^{L-1}\lesssim d_{\mathrm{data}}$. Intuitively, this describes that the network learns to convert the data to a final hidden representation constructed mainly of intrinsic features that are task-relevant, which will be a subset of a larger number of intrinsic features that can be used to describe the data in its original space. Here we will explore this relationship between the intrinsic information content of input data and representations further.
@@
Conclusion

### v4

@@ Introduction
In this work, we first explore how layer-by-layer representation ID progression differs for models trained on datasets beyond ImageNet, including both three other natural image datasets and seven medical image datasets for a wide range of diagnostic tasks. We find that these curves have typically different shapes depending on if the model was trained on natural or medical images. In particular, the depth at which the ID peaks at some $d_{\mathrm{repr}}^{{\mathrm{max}}}$ is noticeably deeper into the model on average for natural image models, which we attribute to the task-relevant features of medical images typically requiring less abstract representations to capture, and so can be “pre-processed” for compression earlier in the network. Next, we explore the relationship between $d_{\mathrm{repr}}^{{\mathrm{max}}}$ and the intrinsic dimension of the raw data in its native space, $d_{\mathrm{data}}$ [27], and discover a new result: there is a strong correlation between $d_{\mathrm{repr}}^{{\mathrm{max}}}$ and $d_{\mathrm{data}}$ across all datasets and models, implying that the maximum intrinsic complexity of learned representations in a model is guided by the intrinsic complexity of the model’s training data.
@@ 1 Experimental Methods
Neural networks work by mapping high-dimensional input data (like images) to significantly lower-dimensional manifolds of learned representations [10, 2] which describe features that generalize to new data, made possible by learning the lower-dimensional manifold structure of such datasets in their native space [7]. The intrinsic dimension of such a manifold can be computed via a relatively simple estimator that utilizes maximum likelihood estimation (MLE) [20, 23] given the high dimensional dataset (in our case, of representations) which we use here following prior work [17, 27] via the implementation of https://github.com/mazurowski-lab/intrinsic-properties. We set the nearest-neighbor hyperparameter to $k=20$ following [17, 16, 27].
@@ 2 Experiments and Results
We hypothesize that the learned representation manifolds of a network may be shaped by the properties and complexity of the dataset manifold in its native (pixel) space. Prior work [17] studied the relationship between the intrinsic dimension of final hidden layer representations for classification networks, $d_{\mathrm{repr}}^{L-1}$, with the intrinsic dimension of the dataset itself, $d_{\mathrm{data}}$, and theoretically and empirically showed that $d_{\mathrm{repr}}^{L-1}\lesssim d_{\mathrm{data}}$. Intuitively, this describes that the network learns to convert the data to a final hidden representation constructed mainly of intrinsic features that are task-relevant, which will be a subset of a larger number of intrinsic features that can be used to describe the data in its original space. Here we will explore this relationship between the intrinsic information content of input data and representations further.
@@
Related Work
Beyond intrinsic dimension (ID), other metrics of the intrinsic complexity, dimensionality, capacity, or related characteristics of a hidden layer’s representations include manifold capacity [3], linear ID [13], local ID [1, 21], and others. We choose to analyze MLE estimator-based ID rather than manifold capacity due to it not requiring specifying labels/a task for the dataset, as manifold capacity focuses on how class representation manifolds are modified through the network, so that our findings may more easily generalize to models for tasks beyond classification. We focus on global nonlinear ID rather than local or linear ID due to its established relationship with important phenomena such as network generalization ability [2]. Other general nonlinear manifold ID estimators beyond MLE include TwoNN [6], GeoMLE [9], and k-NN Graph Distance [11].
Conclusions
@@
We expect that our findings will extend to other tasks beyond binary classification, not just because the ID estimator does not use labels, but also because similar “pre-processing then compression” representation ID progression curves (Fig. 1) have been seen in models for both multi-class classification [2] and other non-supervised, non-vision (as well as non-CNN) models [35]. Future research directions could include (1) further probing our proposed connections between learned feature “abstractness” and peak representation ID, such as by seeing if training for classification tasks with increasingly abstract super-classes (e.g., as in WordNet [25]) results in an increased peak ID depth in the network; as well as (2) attempting to disentangle the effects of dataset ID and feature “abstractness” on peak representation ID using experiments that can control the precise ID of training data, such as in [27].